Define conditional probability with an example?
there are 3 types of conditional probability:
1. the indicate: if antecedent happens, then evidence manifests itself
example -> when tossing a coin, if it lands on tails, then you win the game
2. the subjunctive: if antecedent would happen, then evidence would manifest itself
example -> when a coin was tossed, if it would have landed on tails, i would have won the game
it is recommended for optimal bayesian inferences that your a priori distribution is indicative. If not, you could be dealing with inproper, uninformative, or hyper priors, which make decision-making and posterior determination more complex, if even possible.
Posterior distributions could very well be subjunctive. Suppose i have won the game, i could have tossed tails, but i could also have tossed heads.
What was the motivation for introducing the Gamma distribution?
It can be thought of as a generalization of the Chi-square distribution. See the link to a related WikiAnswer question below.
10 Green marbles, 13 Blue marbles.
Suppose a normal random variable has a mean of 72 inches and a standard deviation of 2 inches. Suppose the random variable X measures the height of adult males in a certain city. One may therefore conclude that approximately 84% of the men in this population are shorter than?
What is the probability of getting a red card from a deck of cards?
Assuming you refer to a normal deck of playing cards, either 0.481 (with jokers) or 0.5, one-half (without).
Describe how the sample size affects the standard error?
Standard error (which is the standard deviation of the distribution of sample means), defined as σ/√n, n being the sample size, decreases as the sample size n increases.
And vice-versa, as the sample size gets smaller, standard error goes up. The law of large numbers applies here, the larger the sample is, the better it will reflect that particular population.
What is the probability that a couple with four children have atleast one girl?
There are 2 ways to do this problem. 1. Go to a Binomial Distribution Table where n = 4 (4 children) and P=0.5(50% probability of a girl). Probability of at least 1 girl = 1 - probability of no girls. From Binomial Distribution Table n = 0 probability is .0625. So, 1 - 0.0625 = .9375 = probability of at least 1 girl. 2. The other way is to list all the possible ways to have 4 children and count the number of ways at least 1 girl exists divided by the total number of ways to have 4 children. There are 42 ways to have 4 children, all 16 listed below: bbbb bbbg bbgb bgbb gbbb bbgg bggb ggbb gbbg gbgb bgbg bggg gggb ggbg gbgg gggg Since 15 of the 16 have at least 1 girl, the Probability of at least 1 girl = 15/16 = 0.9375, the same answer as above.
There is 2 outcomes for flipping the coin, and 6 outcomes for rolling the cube. The total outcomes for both are 2*6 = 12.
If two events are mutually exclusive what is the probability that both occur at the same time?
The probability is 0. Consider the event of tossing a coin . The possible events are occurrence of head and tail. they are mutually exclusive events. Hence the probability of getting both the head and tail in a single trial is 0.
What are the units of measurement for the standard error of mean?
The same units as the mean itself. If the units of the mean, are, for example miles; then the error units are miles.
Can normal distribution be applied on discrete data?
Yes. The normal distribution is used to approximate a binomial distribution when the sample size (n) times the probability of success (p), and the probability of failure (q) are both greater than or equal to 5. The mean of the normal approximation is n*p and the standard deviation is the square root of n*p*q.
How many ways can the word great be arranged?
If you mean re-arranging the letters 'g', 'r', 'e', 'a', and 't' in different sequences, then: We have five letters. The first letter can be any one of the five.
For each choice, the second letter can be any one of the remaining four.
For each choice, the third letter can be any one of the remaining three.
For each choice, the fourth letter can be either one of the remaining two.
Then there is only one letter left to put in the fifth place. So there are (5 x 4 x 3 x 2 x 1) = 120 different ways to carry out this process.
Is there 52 cards in a pack without joker?
Yes there are 52cards in a deck without joker - its all depends on game or country which you play
vijay
What is the ratio of the number of ways an event can occur to the number of possible outcomes?
The term is probability (theoretical probability), or how likely a given event is to occur.
A number cube is rolled 2 times in a row What is the probability of rolling a 3 both times?
When a number cube is rolled twice, there are 36 possible outcomes. (1,1),(1,2),....(6,6).
(3,3) occurs only once. Therefore, the probability of rolling a 3 both times is 1/36.
If you Flip four coins at once what is probability of 2 head and 3 tail?
http://wiki.answers.com/Q/If_you_Flip_four_coins_at_once_what_is_probability_of_2_head_and_3_tail" The probability of flipping four coins and getting 2 heads and 3 tails is ZERO 2 heads and 3 tails requires flipping FIVE coins.
If you flip a coin 3 times how many possible outcomes are there?
There are 2 possibilities for each toss. Since the three tosses are independent (one trial does not affect the outcome of the other trials), there are 2 * 2 * 2 = 8 total possible outcomes.
The outcomes are:
HHH
HHT
HTH
HTT
THH
THT
TTH
TTT
1/3
What is the best way to control events that occur outside your vehicle?
Planning your route is the best way to control events that occur outside your vehicle.
What is the probability of rolling a 7 twice in two rolls of a pair of dice?
The probability of rolling 7 once with two dice is 1 in 6, o 0.1667. The probability, then, of doing that twice in a row is 1 in 36, or 0.02778.